Weekends
Given two dates D1 and D2, write a program to count the number of Saturdays and Sundays from D1 to  D2 (including D1 and D2).
The date in string format is like "8 Feb 2021".
Input

The first line of input will contain date D1 in the string format.
The second line of input will contain date D2 in the string format.
Output

The output should be a single line containing two integers separated by space.
Explanation

For example, if the given dates are "25 Jan 2021" and "14 Feb 2021", the Saturdays and Sundays dates from "25 Jan 2021" to "14 Feb 2021" are 
"30 Jan 2021" is a Saturday
"31 Jan 2021" is a Sunday
"6 Feb 2021" is a Saturday
"7 Feb 2021" is a Sunday
"13 Feb 2021" is a Saturday
"14 Feb 2021" is a Sunday
So the output should be 
Saturday: 3
Sunday: 3
Sample Input 1
25 Jan 2021
14 Feb 2021
Sample Output 1
Saturday: 3
Sunday: 3

Sample Input 2
25 May 2019
22 Dec 2021
Sample Output 2
Saturday: 135
Sunday: 135


1
Expert's answer
2021-05-08T12:32:51-0400

from datetime import date

def dayCount(D1, D2):
    totalDays=(D2 - D1).days
    if(totalDays%7==0):
     print('Total number of saturday and sunday between these two days:',int(totalDays/7))
    else:
     if(totalDays%7<5):
      print('The total number of saturday and sunday between these two days:',int(totalDays/7))
     else:
      if(totalDays%7<6):
       print('Number of saturday:',int(totalDays/7)+1,'and total number of sunday',int(totalDays/7))
    return totalDays

D1 = date(2023, 9, 11)
D2 = date(2022, 1, 9)
print('Total number of days between these two dates',dayCount(D2, D1))

Output: